compute the smallest positive integer x for which (180*x) is a perfect cube




an obtuse triangle has 2 sides of lengths 8 and 11. compute the number of possible integral length of the third side

Answer by Alan3354(69443)

compute the smallest positive integer x for which (180*x) is a perfect cube
180 = 2*2*3*3*5
To be a cube, it needs a 2, a 3 and 2 5's
= 180*2*3*5*5
= 27000 ( = 30^3)
-----------------
an obtuse triangle has 2 sides of lengths 8 and 11. compute the number of possible integral length of the third side
-----------
No side of a triangle can be greater than or equal to the sum of the other 2 sides.
The 3rd side has to be greater than 3 (11-8) and less than 19 (8+11)
19>s>3
-----
oh, integers.
4,5,6,7,8,9,10,11,12,13,14,15,16,17,18
or 4 to 18 --> 15 possible values.
------------------------------
PS It's not 14, 18-4, it's 15.